OpenStudy (anonymous):

Solve x2 + 9x + 9 = 0? Quadratic Formulas

OpenStudy (anonymous):

quadratic formula: x = (-9 +- sqrt(45))/2 = (-9 +- 3*sqrt(5))/2 = 3(-3 +- sqrt(5))/2

OpenStudy (anonymous):

here's a Example Solve the equation: x2 - 9x + 20 = 0 Solution First, factorise the quadratic equation x2- 9x + 20 = 0 Find two numbers which add up to 9 and multiply to give 20. These numbers are 4 and 5. (x - 4) (x - 5) = 0 Now find the value x so that when these brackets are multiplied together the answer is 0. This means either (x - 4) = 0 or (x - 5) = 0 So x = 4 or x = 5. You can check these answers by substitutuing 4 and 5 in to the equation: x2- 9x + 20 Substituting 4 gives: 42 - 9 × 4 + 20 = 16 - 36 + 20 = 0 Substituting 5 gives: 52 - 9 × 5 + 20 = 25 - 45 + 20 = 0 Remember these 3 simple steps and you will be able to solve quadratic equations. quadratic formula: x = (-9 +- sqrt(45))/2 = (-9 +- 3*sqrt(5))/2 = 3(-3 +- sqrt(5))/2

OpenStudy (anonymous):

it will help u more than the solution

OpenStudy (anonymous):

okay.. then what about 2x2 – 8x = –7?

OpenStudy (lgbasallote):

just use what you have learned from their statements...the form of a quadratic equation is ax^2 + bx + c = 0 the quad formula is: x = -b +- sqrt (b^2 - 4(a)(c) ) all over 2a try :DD

OpenStudy (anonymous):

2x^2 - 8x + 7 = 0 (x - 1) x (2x - 7) = 0 You get an F if you can't finish it from here. Late edit: -8x. Crap use the equation to solve. ok. 2X^2 - 8X + 7 = 0 2(X^2 - 4X) + 7 = 0 2(X^2 - 4X + 4) - 8 + 7 = 0 2(X-2)^2 - 1 = 0 (X-2)^2 = 1/2 X - 2 = sqrt(1/2) X = 2 + sqrt(1/2) Or X = sqrt 2 x (sqrt 2 + 1) Or X = (4 + sqrt 2)/2.........your pref.